### v1

Abstract
Ground-based remote sensing cloud image sequence extrapolation is a key research area in the development of photovoltaic power systems. However, existing approaches exhibit several limitations: (1) they primarily rely on static kernels to augment feature information, lacking adaptive mechanisms to extract features at varying resolutions dynamically; (2) temporal guidance is insufficient, leading to suboptimal modeling of long-range spatiotemporal dependencies; and (3) the quadratic computational cost of attention mechanisms is often overlooked, limiting efficiency in practical deployment. To address these challenges, we propose USF-Net, a Unified Spatiotemporal Fusion Network that integrates adaptive large-kernel convolutions and a low-complexity attention mechanism, combining temporal flow information within an encoder-decoder framework. Specifically, the encoder employs three basic layers to extract features. Followed by the unified spatiotemporal module, which comprises: (1) a spatial information branch equipped with a spatial selection module that dynamically captures multi-scale contextual information, and (2) a temporal information branch featuring a temporal agent attention module that effectively models long-range temporal dependencies while maintaining computational efficiency. In addition, a dynamic spatiotemporal module with a temporal guidance module is introduced to enable unified modeling of temporally guided spatiotemporal dependencies. On the decoder side, a dynamic update module is employed to address the common “ghosting effect.” It utilizes the initial temporal state as an attention operator to preserve critical motion signatures. As a key contribution, we also introduce and release the ASI-Cloud Image Sequence (ASI-CIS) dataset, a new large-scale, high-resolution benchmark designed to address the critical limitations of existing public datasets. Extensive experiments on ASI-CIS demonstrate that USF-Net significantly outperforms state-of-the-art methods, establishing a superior balance between prediction accuracy and computational efficiency for ground-based cloud extrapolation. The dataset and source code will be available at https://github.com/she1110/ASI-CIS.
keywords:
[…]
1 Introduction
With the global energy landscape undergoing an accelerated transition toward cleaner sources, photovoltaic (PV) power generation has emerged as one of the fastest-growing sectors in renewable energy due to its advantages of zero-carbon emissions and widespread resource availability [Intro1]. However, the inherent intermittency of solar power, driven by its high dependency on solar irradiance, presents a significant challenge to grid stability. Rapid fluctuations in power output, often caused by fast-moving clouds, can impose substantial pressure on power dispatch systems and energy storage configurations, complicating grid integration [Intro2]. Consequently, the ability to accurately forecast solar irradiance at high temporal resolutions has become a critical enabler for enhancing PV integration capacity, ensuring power system reliability, and facilitating the large-scale deployment of solar technologies.
The output power of PV systems is strongly correlated with solar irradiance, which is primarily modulated by atmospheric factors like cloud cover [Intro3]. Irradiance fluctuations due to cloud obstruction can be broadly categorized into two types: sustained shading (e.g., stratiform cloud coverage) and transient shading (e.g., rapid cumulonimbus movement). While numerical weather prediction (NWP) models can forecast sustained events, transient shading requires real-time, minute-level monitoring to capture its impact [Intro4]. Cloud observation is typically performed using satellite or ground-based remote sensing [Intro5, Intro6]. Satellite-based cloud imagery, with its low spatial resolution (typically >> 1 km) and infrequent updates (typically $\geq$ 30-minute intervals), is inadequate for tracking the localized, dynamic evolution of clouds that cause rapid irradiance changes. In contrast, ground-based cloud imagers provide high-resolution, high-frequency data on cloud structure and dynamics, making them the optimal data source for this task [25TIM]. Therefore, improving the prediction accuracy of ground-based cloud image sequences is critical for achieving precise, ultra-short-term irradiance forecasting, which plays a pivotal role in ensuring the operational stability of grid-connected PV systems [Intro7, Intro8, Intro9].
The task of ground-based remote sensing cloud image sequence extrapolation is fundamentally a deterministic spatiotemporal prediction problem. Advancing the state-of-the-art (SOTA) requires addressing two core technical challenges: a) developing models with the capacity to represent the complex, nonlinear, and multi-scale morphological variations of clouds, and b) achieving this with algorithms that can efficiently capture long-range spatiotemporal dependencies to meet the real-time inference demands of practical applications.
Existing methodologies have sought to address these challenges through two primary avenues: traditional frameworks and, more recently, deep learning-based approaches. Traditional frameworks, which often rely on linear models embedded with vector field representations (e.g., optical flow [Intro10], similarity maximization [Intro11]). However, these methods often suffer from excessive computational overhead and insufficient nonlinear feature representation capabilities, resulting in substantial performance degradation under complex atmospheric conditions.
In response, deep learning techniques have garnered significant attention in spatiotemporal sequence prediction research. Numerous studies have highlighted the effectiveness of recurrent neural networks (RNNs) [Intro12], a classical deep learning algorithm for sequential data processing, in modeling spatiotemporal features. Subsequently, long short-term memory (LSTM) networks [Intro14], as RNN variants, have been successfully applied to capture the temporal dynamics of cloud motion due to their ability to maintain long-range temporal dependencies [Intro13, PredRNN, PredRNN++]. To mitigate the inherent sequential computation constraints of LSTM-based methods, which limit parallelization efficiency, recent works have proposed hybrid architectures integrating convolutional neural networks (CNNs) with LSTM to enhance local spatial feature extraction [ConvLSTM]. Addressing the multi-scale variability inherent in ground-based cloud images, researchers have further developed multi-scale convolutional kernels to extract cloud morphological features across spatial resolutions hierarchically [Intro16, Intro17]. Some innovations like the Motion-Aware Unit (MAU) and various attention mechanisms have been introduced to capture motion patterns and long-range dependencies [MAU, Intro18, Intro19].
While these methods have yielded significant performance gains, critical limitations persist. Firstly, the reliance on static, fixed-size kernels fails to dynamically adapt receptive fields to the diverse and evolving scales of cloud structures, limiting the ability to resolve scale ambiguities in rapidly changing systems. Moreover, the interaction between temporal and spatial feature streams remains underdeveloped, lacking an explicit mechanism for temporal information to guide spatial feature learning. This decoupled approach compromises the capture of long-range relationships. Furthermore, the quadratic complexity of standard self-attention of these solutions is frequently overlooked, becoming a prohibitive bottleneck for the high-resolution imagery and real-time inference required in practical cloud monitoring applications.
The extrapolation of ground-based cloud image sequences is a critical task for enhancing ultra-short-term PV power forecasting. However, existing datasets used in current research suffer from low spatial resolution and hardware-induced visual obstructions, which hinder accurate predictions of PV power generation [MSTANet, CloudPredRNN++]. This limitation primarily stems from the fact that low resolution and hardware obstructions introduce artifact interference, significantly restricting the application of cloud image sequence extrapolation in accurate, real-world PV power forecasting. Consequently, it is particularly important to develop benchmark datasets and techniques for high-resolution, multi-scale, and cross-seasonal cloud image sequence extrapolation.
To address these distinct challenges, this article introduces the Unified Spatiotemporal Fusion Network (USF-Net), a novel architecture that integrates adaptive large-kernel convolutions with a low-complexity attention mechanism within a unified encoder-decoder framework. Specifically, the encoder employs depthwise separable (DW) convolutions and squeeze-excitation (SE) blocks for hierarchical downsampling and multi-scale feature extraction. Following the encoder, we propose a unified spatiotemporal module (USTM) comprising three core components: 1) a spatial information branch, in which a spatial selection module (SSM) is designed to extract multi-scale cloud context through adaptive receptive field adjustment dynamically; 2) a temporal information branch incorporating a temporal agent attention module (TAM) is introduced to capture long-range temporal dependencies in cloud sequences while reducing computational complexity; and 3) a dynamic spatiotemporal module: a temporal guidance module (TGM) fuses spatial and temporal features, enabling unified modeling of temporal flow-guided spatiotemporal dependencies. For the decoder part, a dynamic update module (DUM) is implemented to mitigate the “ghosting effect” by contextual information decay during upsampling. In the DUM, the initial temporal state is employed as gating units to reweight spatiotemporal features, refining multi-scale representations while preserving critical cloud motion signatures. Finally, we publicly develop and release a novel ground-based cloud image sequence dataset, comprising large-scale, high-resolution sequences captured under diverse meteorological conditions. The main contributions of this article are summarized as follows.
  1. 1.
    A novel unified spatiotemporal architecture, USF-Net, is proposed. It explicitly integrates temporal flow information to guide spatial feature learning, which enhances the coherence of temporal-spatial dependencies modeling and significantly improves prediction accuracy for complex cloud dynamics.
  2. 2.
    A Unified Spatiotemporal Module (USTM) is designed to serve as the core of the network. It features a Spatial Selection Module (SSM) for dynamic, adaptive multi-scale feature extraction and a low-complexity Temporal Agent Attention Module (TAM) that effectively balances predictive accuracy with computational efficiency.
  3. 3.
    A Dynamic Update Module (DUM) is introduced in the decoder. It leverages initial temporal states as an attention operator to reweight upsampled features, preserving critical motion signatures and effectively mitigating the “ghosting effect.”
  4. 4.
    The introduction and public release of the ASI-Cloud Image Sequence (ASI-CIS) dataset. As a major contribution to the community, ASI-CIS is a newly introduced, large-scale, high-resolution, multi-seasonal benchmark that addresses key limitations of previous datasets. It offers a more realistic and challenging foundation for advancing ground-based cloud extrapolation models. Extensive experiments on ASI-CIS show that USF-Net outperforms state-of-the-art methods in both prediction accuracy and computational efficiency.
The rest of this paper is organized as follows. In Section 2, we introduce the related works of ground-based cloud image sequence precidtion. In Section 3, we introduce the detailed structure of the proposed method. In Section 4, we evaluate the performance of our proposed methods. Finally, Section 5 concludes the paper.
2 Related Works
Accurate prediction of cloud image sequence extrapolation plays a pivotal role in enhancing the operational stability of grid-connected PV systems. Current methodologies are broadly divided into two categories: traditional methods and deep learning-based methods.
2.1 Traditional methods for cloud image sequence extrapolation
Conventional approaches predominantly employ optical flow (OF) algorithms for cloud image sequence extrapolation. OF is utilized to estimate the instantaneous velocity field of pixel-wise motion in ground-based cloud images, capturing spatiotemporal trends of cloud dynamics. Several studies have adopted OF-based methods for ground-based remote sensing cloud image sequence extrapolation. Omnidirectional optical flow tracking frameworks are proposed to establish quantitative relationships between cloud motion directionality and velocity magnitudes in cloud dynamics [Re1, Re2]. Boundary information within cloud imagery constitutes a critical determinant of prediction accuracy. Chang et al. implemented the Horn-Schunck OF algorithm to compute velocity variations for each pixel, augmenting motion field estimation through supplementary information integration [Re3]. However, these methods incur substantial computational overhead. Conversely, Wang et al. introduced a mathematical analysis framework to characterize inter-frame disparities, achieving reduced computational resource consumption while maintaining predictive performance [Re4].
Despite the operational feasibility of the aforementioned methodologies in executing cloud image sequence extrapolation tasks, conventional approaches exhibit persistent limitations, including being computationally prohibitive and exhibiting limited motion modeling capabilities. These deficiencies manifest in their inability to capture temporal motion patterns under complex cloud regimes characterized effectively. The rapid advancement of DL techniques in computer vision has spurred transformative progress in this domain. Numerous studies have applied DL to the prediction of cloud image sequence extrapolation, achieving remarkable advancements.
2.2 DL-based methods for cloud image sequence extrapolation
In the historical development of DL, recurrent neural networks and their variant long short-term memory networks have served as pioneering methods for temporal sequence prediction. Several studies have successfully employed LSTM architecture to capture long-range dependencies in spatiotemporal sequences. However, standalone LSTM models exhibit elevated computational costs while struggling to extract complex spatial information effectively. Consequently, numerous works integrate CNN with LSTM frameworks to jointly extract localized spatial features and temporal information for sequential motion prediction. PredRNN enhanced the ConvLSTM architecture by introducing a zigzag memory flow mechanism to model short-term spatiotemporal dynamics [PredRNN]. Subsequently, PredRNN++ incorporated a gradient highway unit (GHU) to mitigate the gradient vanishing issue prevalent in LSTM-based models while integrating Causal-LSTM modules to strengthen spatial feature representation and short-term temporal modeling [PredRNN++]. Li et al. proposed cascaded Causal-LSTM layers to improve short-term prediction accuracy for cloud imagery. The model was augmented by GHUs with auxiliary skip connections to enhance spatiotemporal uniformity in modeling [CloudPredRNN++]. Nevertheless, ground-based remote sensing cloud image exhibits high-resolution cloud formations with multi-scale variations under complex meteorological conditions. Existing methodologies remain suboptimal for cloud sequence extrapolation tasks. To capture the dynamic states of cloud clusters at varying scales within cloud imagery, several studies have integrated multi-scale convolutional kernels. For instance, MSSTNet employs 3D convolutions with diverse kernel sizes to enhance the capacity of the model for multi-resolution image forecasting [Re5]. Wang et al. introduced 3D tensor augmentations within LSTM architectures to expand the effective receptive field [Re6]. However, the adoption of 3D convolution operations incurs significant computational overhead. The MSTANet employs multi-scale large kernels to aggregate multi-scale contextual information from cloud imagery while leveraging depthwise separable convolutions to construct large kernels with reduced computational complexity [MSTANet]. Accurate cloud image sequence extrapolation serves as a critical factor in ultra-short-term PV power forecasting, and the modeling of long-range spatiotemporal dependencies becomes particularly paramount. To this end, Chang et al. designed an MAU that simultaneously enlarges the model’s receptive field and captures spatial motion patterns across cloud sequences [MAU]. The Motion RNN introduces a Motion RGU module to unify transient variation modeling and motion trend representation [Re7]. When embedded within RNN architectures, this approach significantly improves spatiotemporal prediction accuracy under complex meteorological scenarios.
Attention mechanisms have been demonstrated as an effective approach for establishing long-range dependencies, facilitating the extraction of temporal features in sequence prediction tasks. The STANet introduces a context gating unit (CGU) as an attention mechanism to unify the modeling of instantaneous cloud characteristics and motion trends [STANet]. Similarly, SimvPV2 incorporates a gated spatiotemporal attention module to enforce spatiotemporal consistency in sequence modeling [Re8]. Tan et al. decomposed temporal attention into intra-frame static attention and inter-frame dynamic attention through a dedicated temporal attention unit, capturing spatial features and temporal correlations [TAU]. Li et al. integrated a self-attention memory unit into the Cascaded Causal LSTM (CCLSTM) framework to extract long-term dependencies, enabling spatiotemporal modeling for cloud sequence prediction [CCLSTM]. Furthermore, the MSTANet introduces a multi-scale temporal attention mechanism that combines local temporal variations and global temporal variations, significantly enhancing the network’s temporal modeling capacity for cloud image extrapolation tasks [MSTANet].
Unfortunately, while the methods above demonstrate commendable performance in cloud image sequence extrapolation tasks, they encounter three notable limitations. First, these approaches solely employ multi-scale convolutional kernels to capture contextual information, lacking adaptive mechanisms to extract features at varying resolutions dynamically. In addition, during spatiotemporal dependency modeling, the absence of temporal guidance hinders the unified integration of spatial and temporal information, resulting in suboptimal long-term dependency capture. Furthermore, existing attention mechanisms for temporal feature extraction neglect to balance computational complexity with prediction accuracy, leading to inefficiencies in practical deployment.
3 Proposed Method
This section provides a detailed exposition of the USF-Net architecture, beginning with the problem formulation and a high-level overview, followed by in-depth descriptions of its novel components: the Unified Spatiotemporal Module (USTM), the Dynamic Update Module (DUM), and the composite loss function.
Figure 1: (a) illustrates multi-scale cloud movement. The red, yellow, and blue blocks represent displacement vectors of large, medium, and small-scale clouds, respectively. The arrow indicates the direction of the movement trend. (b) demonstrates “ghosting effects” in cloud image sequence extrapolation. The orange block denotes ground truth (GT), whereas the green block indicates extrapolation results.
1
Input : Input cloud sequence $X_{t}^{T}=\{x_{i}\}_{t+1}^{T}$
Output : Extrapolated sequence $Y_{T+1}^{T+\tau}=\{y_{i}\}_{T+1}^{T+\tau}$
2
31ex
4repeat
5   Let layer $i$ = 1, lossloss = 0.0
6 $L=\lambda_{1}L_{M}+\lambda_{2}L_{MS}+\lambda_{3}L_{C}$ ($L$: loss function)
7 for $i\leftarrow 1$ to 33 do
    $X_{B}\leftarrow N_{i}(X_{t}^{T})$ ;
    // Encoder layer $i$
8    if $i=3$ then
9       $X_{T}\leftarrow X_{B}$
10      end if
11
12   end for
13
14 $X_{S}\leftarrow\text{SiB}(X_{B})$
15 $X_{T}\leftarrow\text{TiB}(X_{T})$
16 $X_{D}\leftarrow\text{DSM}(X_{S},X_{T})$
17 $D_{4}\leftarrow\text{DUM}(X_{D},X_{T_{0}})$
18 for $k\leftarrow 3$ to 11 do
    $D_{k}\leftarrow\text{UP}(D_{k+1})$ ;
    // Decoder layer $k$
19
20   end for
21 $Y\leftarrow\text{Conv}_{1\times 1}(D_{1})$
22 $\mathcal{L}\leftarrow\lambda_{1}L_{M}+\lambda_{2}L_{MS}+\lambda_{3}L_{C}$
 $\theta\leftarrow\theta-\nabla_{\theta}\mathcal{L}$ ;
 // Parameter update
23
24until convergence;
return $Y_{T+1}^{T+\tau}$
Algorithm 1 Procedure of the USF-Net
3.1 Formulation and Architectural Overview
The ground-based cloud image sequence extrapolation task is formulated as follows: given an input sequence of $T$ historical frames, denoted as $X_{t}^{T}=\left\{x_{i}\right\}_{t+1}^{T}$, where each frame $x_{i}\in\mathbb{R}^{C\times H\times W}$, the objective is to predict a sequence of $\tau$ future frames after $T$, $Y_{T+1}^{T+\tau}=\left\{y_{i}\right\}_{T+1}^{T+\tau}$, where $y_{i}\in\mathbb{R}^{C\times H\times W}$. The model, parameterized by $\theta$, learns a mapping $F_{\theta}:X_{t}^{T}\rightarrow Y_{T+1}^{T+\tau}$ by maximizing the log-likelihood of the predicted frames (e.g., cloud images) with respect to the ground-truth counterparts.
Specifically, $C$, $H$, and $W$ correspond to the channel, height, and width dimensions of the images, respectively. Notably, the input and output of cloud extrapolation tasks are structured as tensors, i.e., $X_{t}^{T}\in\mathbb{R}^{T\times C\times H\times W}$ and $Y_{T+1}^{T+\tau}\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ denotes the frame rate. Formally, our objective prediction optimization model can be parameterized as $\theta$:
$\theta_{T}=arg\underset{\theta}{max}{\textstyle\sum_{i=T+1}^{T+\tau}}logP\left(y_{i}\mid x_{i};\theta\right)$ (1)
where the predictive model constitutes a learnable mapping $\theta_{T}$ that maximizes the log-likelihood between predicted cloud images and the ground-truth counterparts.
Existing methods have demonstrated that the ground-based cloud image sequence extrapolation task faces several critical challenges. As illustrated in Fig. 1 (a), the scale-variant properties of cloud formations during motion introduce inaccuracies in sequence extrapolation due to multi-scale variations. Moreover, Fig. 1 (b) reveals that prevalent approaches suffer from partial contextual information loss during the decoder phase, leading to the emergence of “ghosting effects” that complicate cloud motion trajectory prediction. These motivate the design of a multi-scale network model with a spatiotemporally unified architecture, aiming to improve the precision of cloud sequence prediction while simultaneously enhancing inference efficiency.
Figure 2: (a) The structure of the proposed USF-Net is composed of three parts: the encoder comprises three Basic Layers, the USTM and the decoder comprises a dynamic update module (DUM). $C_{i}$ denotes the channel of the feature map. (b) The structure of the encoder, where $N_{1}$, $N_{2}$, and $N_{3}$ are 2, 2, and 3, respectively. The output of the encoder is $X_{B}$. (c) The specific structure of the Basic Layer. (d) The diagram of the proposed Unified SpatioTemporal Module (USTM) comprises three core components: a spatial information branch (SiB), a temporal information branch (TiB), and a dynamic spatiotemporal module (DSM). The output of the USTM is $X_{T}$.
3.2 Overview of Structure
The proposed USF-Net, as shown in Fig. 2 (a), adopts a universal encoder-decoder architecture. The procedure of our USF-Net is described in Algorithm 1. A unified spatiotemporal module is introduced at the bottleneck to enhance the accuracy of ground-based cloud image sequence extrapolation by explicitly incorporating temporal guidance. As shown in Fig. 2 (a) and (b), in the encoder part, the input $X_{t}^{T}\in\mathbb{R}^{B\times T\times C\times H\times W}$ is first processed by three basic layers. Each layer consists of a $3\times 3$ DW convolutional layer for local feature extraction, followed by layer normalization, a DW convolutional layer, a Squeeze-and-Excitation (SE) block and a convolutional layer. The residual connection is used to enrich the representation capabilities. We denote the feature map of each layer as $f_{i}(i\in[1,2,3])$. The size of $f_{i}$ is $T\times 2^{i-1}\cdot 64\cdot\left(H\times W\right)/2^{i+1}(i\in[1,2,3])$, ultimately generating an intermediate feature map $X_{B}\in T\cdot C_{3}\cdot\left(H\times W\right)/16$. This feature map $X_{B}$ is then fed into the USTM, which comprises dual spatial and temporal branches coupled with a DSM. The spatial branch is employed to extract multi-scale spatial information, while the temporal branch is employed to capture temporal dependencies from the input sequence. These temporal cues are subsequently channeled through the DSM to guide spatial feature fusion, achieving unified spatiotemporal modeling. In the decoder part, to mitigate the “ghosting effect” caused by contextual information loss, a novel gated unit termed the DUM is designed to enable temporally guided spatial feature refinement. Within DUM, initial temporal features are reweighted via a gate mechanism to refine multi-scale spatiotemporal representations, preserving contextual coherence. It is noteworthy that USF-Net jointly extracts spatial features and temporal information, leveraging temporal flows to guide the generation of high-precision semantic representations. Consequently, our architecture can be interpreted as a spatiotemporally unified framework optimized for ground-based cloud image extrapolation, balancing computational efficiency with prediction accuracy. The details of our method are as follows.
Figure 3: The structure of the proposed SiB. The SSM employs explicitly decomposed convolution operations to generate varying receptive field sizes, thereby enhancing the network’s multi-scale representational capacity.
3.3 Unified SpatioTemporal Module
Learning multi-scale contextual information from scale-variant cloud imagery is critical for ground-based cloud image extrapolation tasks, as it facilitates precise spatial feature extraction. In real-world scenarios, complex meteorological conditions, such as variable wind speeds and diverse atmospheric flow patterns, result in non-stationary and nonlinear motion dynamics within cloud sequences. Therefore, establishing effective spatiotemporal dependencies is essential to achieve accurate predictions in cloud image extrapolation tasks.
However, most existing cloud image extrapolation methods that apply large-scale kernels ignore the importance of dynamic adaptive selection in capturing multi-scale fine-grained features from cloud imagery. In addition, spatiotemporal feature extraction in these approaches is often decoupled, lacking interactive guidance between spatial and temporal representations, and ignoring the influence of temporal information flows on spatial feature learning. Furthermore, prevalent methods employing self-attention mechanisms or LSTM-based recurrent architectures for temporal modeling incur prohibitive computational complexity due to their quadratic or sequential operational paradigms, limiting practical applicability in resource-constrained scenarios.
To this end, we design a unified spatiotemporal module to improve the ability to explore multi-scale contextual information of cloud imagery adaptively. Temporal dependencies are efficiently captured using a low-complexity attention mechanism. By incorporating TGM to inject temporal flows into spatial representations, the model ensures holistic spatiotemporal dependency modeling, significantly improving the precision of ground-based remote sensing cloud image extrapolation.
[…]
3.3.1 Spatial Information Branch
Cloud formations in natural meteorological conditions often exhibit complex nonlinear and non-stationary motion, leading to multi-scale variations in cloud image extrapolation tasks. To achieve precise predictions for ground-based remote sensing cloud image extrapolation, comprehensive and dynamically adaptive multi-scale contextual information is essential.
To this end, a spatial information branch with SSM is proposed to enhance the capability of dynamically extracting multi-scale contextual features, as shown in Fig. 3. Inspired by the Large Kernel Selection (LSK) mechanism [LSK], the dynamic adaptive large convolutional kernel adaptively adjusts the receptive field for each spatial target based on the scale of cloud formations. Specifically, the SSM is passing through a DW convolutional layer and a GeLU activation function to balance feature extraction efficiency with nonlinear representation capacity. To enhance the ability of the network to focus on the most relevant spatial contextual regions for multi-scale cloud formation extraction, multi-scale convolutional kernels are employed to adaptively select spatial features, where each kernel $K_{i}$ generates a corresponding feature map $X_{K_{i}}$. This process can be formalized as:
$X_{K_{i}}=DW\left(K_{i}\right)$ (2)
where $i$ denotes the number of large convolutional kernels. In this paper, kernels with configurations $K_{1}=3$, $d_{1}=1$ (dilation rate) and $K_{2}=7$, $d_{2}=3$ are utilized. As demonstrated by the following formula:
$RF_{1}=k_{1}$ (3)
$RF_{2}=d_{2}\left(k_{2}-1\right)+RF_{1}$ (4)
the proposed selection module effectively performs an explicit decomposition of a large kernel ($K=21$). By progressively increasing the kernel size and dilation rate, the explicitly decomposed convolution operations generate varying receptive field sizes, enhancing the network’s multi-scale representational capacity while significantly reducing parameter count.
Subsequently, the multi-scale feature maps are fused to generate $X_{K}$, which encapsulates diverse receptive fields. To enable interaction between cross-spatial features, channel-wise average pooling and max pooling are applied to $X_{K}$, followed by a convolutional layer (Conv) to produce the spatial interaction attention map $\widetilde{SA}$. This process is mathematically formalized as:
$X_{K}=Cat\left(X_{K_{1}},X_{K_{2}}\right)$ (5)
$\widetilde{SA}=Conv\left(Avg\left(X_{K}\right),Max\left(X_{K}\right)\right)$ (6)
where $Cat$ denotes the concatenation of the channel, $Avg$ and $Max$ denote the average pooling and max pooling, respectively.
To achieve dynamic adaptive extraction of multi-scale cloud formations, the attention operators $\widehat{SA}_{i},i\in\left[1,2\right]$ for distinct receptive fields are derived by applying a sigmoid function to $\widetilde{SA}$. The feature maps from the decomposed large-kernel sequence are weighted by their corresponding selection weights, and the fused features are processed via a Conv layer to obtain the multi-scale spatial selection attention map $X_{SA}$. Finally, the spatial branch’s output $X_{S}$ is computed by combining $X_{B}$ with $X_{SA}$, as expressed in:
$\widehat{SA}_{i}=\sigma\left(\widetilde{SA}_{i}\right)$ (7)
$X_{SA}=Conv\left(Cat\left(\widehat{SA}_{1}\ast X_{K_{1}}\widehat{SA}_{2}\ast X_{K_{2}}\right)\right)$ (8)
$X_{S}=X_{SA}\otimes X_{B}$ (9)
where $\otimes$ denotes the element-wise production. By dynamically adjusting the receptive fields of spatial targets within the spatial branch, the proposed method effectively captures contextual information across varying cloud scales.
3.3.2 Temporal Information Branch
Cloud exhibits non-stationary motion over time due to meteorological and temporal influences, necessitating the acquisition of temporal information from cloud imagery sequences to model their motion trends accurately. Conventional approaches often employ recurrent networks or attention mechanisms to extract temporal dependencies from sequences. However, recurrent networks suffer from an inability to parallelize data processing, impotent the rendering of real-time, precise prediction due to the prohibitive computational costs. Self-attention mechanisms in most existing methods introduce quadratic complexity, resulting in substantial computational overhead. Therefore, inspired by agent attention, we propose a TAM, which synergizes the high precision of Softmax attention and the low complexity of Linear attention, achieving a favorable trade-off between computational efficiency and representational capacity for spatiotemporal modeling.
As shown in Fig. 4 (a), the proposed temporal information branch comprises two stacked components: conv embedding (CE) and TAMs. Different from conventional Vision Transformers (ViTs) that partition images into non-overlapping patches via linear projections, we select a CNN layer that replaces the patch embedding (PE) layer, as this design choice mitigates spatial information degradation and preserves multi-scale contextual coherence in cloud imagery. This method inherently retains spatial and positional information without requiring auxiliary positional encoding. The input channel dimension aligns with the embedding dimension in ViT architectures. Specifically, as shown in Fig. 4 (b), a CBR layer, consisting of a $3\times 3$ convolution, batch normalization (BN), and ReLU activation, is employed to extract spatial information. A DW convolutional layer is then integrated with residual connections to reinforce prior knowledge within the embedding process.
Figure 4: (a) The overall structure of the proposed TiB. (b) The proposed CE consists of a $3\times 3$ convolution, batch normalization (BN), ReLU activation, and a DW convolutional layer with residual connections. (c) The proposed DSM, the dashed line denotes the feature flow of Agent attention, and the solid line denotes the feature flow of Softmax attention.
Most existing ViT-based approaches for extracting temporal dependencies via softmax attention often incur excessive computational complexity. While linear attention reduces computational overhead, it compromises model expressiveness, leading to inadequate representation capacity of the network. To this end, a novel attention mechanism, TAM, is proposed, integrating the advantages of both paradigms, enabling effective temporal dependency extraction in cloud imagery sequences as shown in Fig. 4 (c).
Different from the original agent attention, the proposed method does not employ standard projection matrices to derive the query ($Q$) and key ($K$) in the attention mechanism. Instead, customized convolutional kernels are constructed to establish independent adaptive receptive fields for each pixel in cloud imagery, thereby modeling long-range dependencies. These processes can be formally expressed as:
$X_{Q_{ij}}=\sum_{l=-1}^{1}\sum_{g=-1}^{1}E_{2+l,2+g}^{q}X_{i+l,j+g}$ (10)
$X_{K_{ij}}=\sum_{l=-1}^{1}\sum_{g=-1}^{1}E_{2+l,2+g}^{k}X_{i+l,j+g}$ (11)
where $E^{q}$ and $E^{k}\in\mathbb{R}^{T\times C\times 3\times 3}$ denotes a learnable projection matrix that aggregates features from the $3\times 3$ neighborhood of adjacent pixels into $X_{i,j}\in\mathbb{R}^{T\times C\times N}$. While analogous to the linear projection in ViT, this approach generates adaptively learned receptive fields, enhancing the capability of the network to perceive multi-scale variations in cloud imagery through dynamic spatial adaptation.
Subsequently, an agent token, $X_{A}\in\mathbb{R}^{T\times C\times n},n\ll N$, is generated via pooling operations. Notably, the tokens $n$ are set as a small hyper-parameter to achieve a linear computation complexity while maintaining global context modeling capability. Specifically, $X_{A}$ is first treated as the $Q$ to compute attention scores with $X_{K}$ and $X_{V}$, yielding the agent feature $A_{F}$. Then, $A_{F}$ is utilized as the $V$, while $X_{A}$ serves as the $K$, to perform a second attention computation with $X_{Q}$, producing the final attention feature $A$. These processes are formalized as:
$X_{A}=Pooling\left(X_{Q}\right)$ (12)
$A_{F}=Softmax\left(X_{A},\left(X_{K}\right)^{T}\right)X_{V}$ (13)
$A=Softmax\left(X_{Q},\left(X_{A}\right)^{T}\right)A_{F}$ (14)
where $Softmax$ denotes softmax attention function. In this way, we avoid expensive computational costs while preserving the information interaction between $Q$ and $K$. The feature $A$ is then processed through a DW convolutional layer combined with residual connections to obtain diversity in temporal feature information. Finally, following a DW convolution and BN, the output $X_{T}$ of the temporal branch is obtained. This process is expressed as:
$X_{T}=CB\left(Cat\left(DW\left(X_{V},A\right)\right)\right)$ (15)
Figure 5: The structure of the proposed DSM. The bottom-hand side of the figure shows the structure of the TGM in detail. The learnable dynamic convolution kernels are generated by applying weighted guidance from temporal flow information to spatial feature maps utilizing temporal flow information.
3.3.3 Dynamic Spatiotemporal Module
To ensure robust spatiotemporal dependency modeling in cloud image sequence extrapolation, we design a dynamic temporal module with TGM that applies weighted guidance from temporal flow information to spatial feature maps. Temporal dynamics enhance the capacity of the network to capture multi-scale contextual features and model long-term dependency.
[…]
Figure 6: The structure of the proposed DUM. The $X_{e}$ and $X_{M}$ on the left side of the figure are the temporally guided spatiotemporal feature and the initial temporal information from TAM, respectively. The $W$ and $\sigma$ on the right side represent the matrix and active function in the gate unit, respectively.
3.4 Dynamic Update Decoding
The “ghosting effect” is a prevalent challenge in ground-based remote sensing cloud image sequence extrapolation tasks. The decoder of most existing methods is designed to fuse the same scale feature maps via a lateral connection. However, these methods lack temporal information guidance between hierarchical features and ignore the global feature correlation between the local information of different layers.
To this end, we propose a decoder with a DUM to enhance the ability of the method to capture global temporal flow correlations across different layers, which obtains the relationship of different level features with long-range dependencies. Specifically, we build a gate unit with initial temporal information $X_{T_{0}}$ to prevent the temporally guided spatiotemporal context $X_{D}$, derived from the USTM, from being diluted. As shown in Fig. 6, our TGM samples $X_{D}$ and $X_{T_{0}}$ for gate operation. One branch adopts the $X_{T_{0}}$ as the attention operator of the attention mechanism. The other adjusts the channel information adaptively by re-weighting the $X_{D}$ in the decoder through initial temporal flow information. Then, the features of the two branches are fused by a pointwise (PW) convolution to refine the fused information. The global features $D_{4}$ with temporal flow information are generated through the dynamic update decoding, which is conducive to obtaining the long-range dependence of image sequences. The $D_{4}$ will restore the feature scale as the third layer by passing the upsampling operation to obtain the third feature map. The similar processing of the next layer in the decoder will be repeated, and we can obtain the fused feature $D_{i}$. The entire process can be mathematically formulated as follows:
$D_{i}=\begin{cases}\mathrm{DUM}\left(X_{D},X_{T_{0}}\right)&\text{ if }i=4\\ \mathrm{UP}\left(D_{i+1}\right)&\text{ if }i=1,2,3\end{cases}$ (16)
$\mathrm{DUM}=Cat\left(X_{D},\mathrm{PWconv}\left(\mathrm{gate}\left({X_{T}}_{0},X_{D}\right)\right)\right)$ (17)
$\mathrm{gate}\left(X_{D},X_{T_{0}}\right)=\left(\omega_{1}\cdot X_{D}+b\right)\otimes\sigma\left(\omega_{2}\cdot X_{T_{0}}+c\right)$ (18)
where $\mathrm{PWconv}$ denotes the point-wise convolution, which is used to aggregate the features, $\mathrm{gate}$ denotes the gate unit. $\omega$ denotes the learning matrix, $b,c$ denotes the bias term, $\otimes$ denotes element-wise production, and $\sigma$ denotes the sigmod function. The $\mathrm{UP}$ operation is composed of two convolutional operations (the kernel size is 3) and bilinear interpolation with the scale factor is set to 2.
The final feature map of the last layer undergoes a convolution to restore the same size as the original input images. Then, a $1\times 1$ convolution is used to adjust the number of channels to make the final prediction.
3.5 Loss Function
The selection of appropriate loss functions plays a critical role in enhancing the robustness of the network. To improve the prediction accuracy of cloud image sequences, the mean squared error (MSE) loss is adopted based on the community-related works [MSTANet, STANet] to evaluate the global correlation between the ground truth (GT) $y_{i}\in\mathbb{R}^{T\times C\times H\times W}$ and predicted results $y\in\mathbb{R}^{T\times C\times H\times W}$. The formulation of the MSE loss $L_{M}$ is as follows:
$L_{M}=\frac{1}{N}\sum_{i=1}^{N}\left(\widehat{y_{i}}-y_{i}\right)^{2}$ (19)
where $N$ denotes the total number of samples.
However, the MSE loss function overlooks local structural features, which can lead to significant deviations and semantic information loss in cloud image sequence extrapolation tasks. To address this limitation, we introduce the multi-scale structural similarity (MS-SSIM) loss function to preserve edge details and structural information. The MS-SSIM is an enhanced version of the Structural Similarity Index (SSIM), incorporating structural similarity optimization across varied resolution levels. By improving robustness to scale variations in the target, MS-SSIM is particularly suitable for cloud image sequence extrapolation tasks characterized by scale-varying cloud formations.
Firstly, an $S$-level Gaussian pyramid downsampling is performed on the images $y_{i}$ and $\widehat{y_{i}}$, generating multi-scale image pairs $\left\{y_{j},\widehat{y_{j}}\right\}_{j=1}^{L}$ (empirically set as $L$ = 5). Three SSIM components, including luminance $l_{j}$, contrast $c_{j}$, and structure $s_{j}$ as follows:
$l_{j}(y,\hat{y})=\frac{2\mu_{y}\mu_{\hat{y}}+C_{1}}{\mu_{y}^{2}+\mu_{\hat{y}}^{2}+C_{1}}$ (20)
$c_{j}(y,\hat{y})=\frac{2\sigma_{y}\sigma_{\hat{y}}+C_{2}}{\sigma_{y}^{2}+\sigma_{\hat{y}}^{2}+C_{2}}$ (21)
$s_{j}(y,\hat{y})=\frac{\sigma_{y\hat{y}}+C_{3}}{\sigma_{y}\sigma_{\hat{y}}+C_{3}}$ (22)
where $\mu$, $\sigma$, and $\sigma_{y\hat{y}}$ denote mean, standard deviation, and covariance, respectively. $C_{1}$, $C_{2}$, and $C_{3}$ are all constants. Then, the MS-SSIM value is derived by weighted aggregation of the SSIM components across all scales:
$\operatorname{MS-SSIM}=\left[l_{j}^{\alpha}\cdot\prod_{j=1}^{L}c_{j}^{\beta}s_{j}^{\gamma}\right]$ (23)
where $\alpha$, $\beta$, and $\gamma$ are empirically determined weighting exponents ($\alpha$=1, $\gamma$$\beta$=0.0448 by convention). The MS-SSIM loss $L_{MS}$ is defined as: $L_{MS}$=1-MS-SSIM.
Additionally, to emphasize the weight of the first frames in the predicted sequence, the cross-entropy (CE) loss is augmented with a weighting factor $\tau$ (empirically set to 0.9 in this work), formulated as:
$L_{C}=\sum_{i=1}^{T}\tau^{i}{L_{CE}}^{t+i}$ (24)
where $t+i$ denote the future timestep, $L_{CE}$ denote the CE loss. Finally, the loss of our USF-Net can be formulated as:
$L=\lambda_{1}L_{M}+\lambda_{2}L_{MS}+\lambda_{3}L_{C}$ (25)
where $\lambda_{1}=0.7$, $\lambda_{2}=0.2$, and $\lambda_{3}=0.1$.
4 Experimental Results and Discussions
This section details the comprehensive experimental validation of USF-Net. We describe the newly created ASI-CIS dataset, the evaluation metrics, and implementation details. We then present a rigorous comparative analysis against SOTA methods, followed by extensive ablation studies to verify the contribution of each novel component.
Table 1: ASI-CIS Dataset Summary.
Weather Size Number / Sequences
Sunny 512 $\times$ 512 28,420 / 1421
Cloudy/Rainy 512 $\times$ 512 11,580 / 579
4.1 Dataset
A significant barrier to progress in ground-based cloud extrapolation has been the lack of high-quality, large-scale public datasets. The performance of deep learning models is fundamentally tied to the data they are trained on, yet existing benchmarks suffer from critical limitations that fail to capture the true complexity of cloud dynamics. For instance, the popular TSISD dataset [STANet] features a low spatial resolution of only (224 $\times$ 224 $\times$ 3 for each image) and contains significant visual occlusions from camera hardware, which can introduce artifacts and degrade prediction precision.
To address this gap and provide a more challenging and realistic benchmark for the research community, we developed and publicly release the ASI-Cloud Image Sequence (ASI-CIS) dataset, a core contribution of this paper. In this dataset, 1, 400 sequences are used for training and 700 sequences are used for testing. To eliminate data similarity between splits, the training and testing sets are acquired from temporally distinct periods, with validation performed via five-fold cross-validation. The size of each image is 512 $\times$ 512 pixels, and consecutive frames are captured at 30-second intervals using the All Sky Smage (ASI-DC-TK02). This device is located in Xiqing District, Tianjin, China (geographic coordinates: 117.03°E, 39.10°N). Data collection spans multiple seasons, with acquisition times ranging from 08:00 to 17:00 local time, ensuring robust coverage of diverse meteorological conditions for ground-based cloud image sequence extrapolation tasks. Table 1 summarizes the ASI-CIS dataset according to various weather conditions and quantities. The dataset includes samples under diverse weather conditions, including sunny, cloudy/rainy scenarios, with 1421 and 579 sequences, respectively.
Figure 7: (a) and (b) display valid acquisition samples under sunny and cloudy/rainy conditions, respectively; (c) displays samples rendered unsuitable for sequence extrapolation tasks due to complex cloud configurations encountered during adverse meteorological conditions such as precipitation events.
The observed data imbalance primarily stems from the temperate monsoon climate in the acquisition region, where clear-sky conditions occur significantly more frequently than cloudy/rainy weather. This inherent meteorological distribution directly induces substantial disparities in raw data collection volumes. Furthermore, suboptimal acquisition conditions during precipitation events, which are characterized by heavy rainfall, low illumination, and potential lens contamination, frequently yield poor-quality imagery. Cloudy conditions present additional challenges due to complex cloud structures (e.g., stratocumulus-cumulus mixtures) that exceed preset sensor parameters, rendering cloud formations unrecognizable. Representative samples across weather conditions are shown in Fig. 7.
4.2 Evaluation Metrics
Round-based remote sensing cloud image sequence extrapolation task is essentially an image tracking task. Therefore, to evaluate the performance of our proposed method, several common metrics are employed, including the Mean Squared Error (MSE), Structural Similarity Index (SSIM), and Peak Signal-to-Noise Ratio (PSNR).
The MSE indicates the average pixel-wise discrepancy between the predicted result and GT by computing the squared difference across corresponding pixels. The specific formulations are as follows:
$\mathrm{MSE}=\frac{1}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}(I(i,j)-K(i,j))^{2}$ (26)
where $I$ and $K$ denote the predicted and GT images, respectively, and $m\times n$ denotes the image dimensions.
SSIM assesses visual quality by comparing luminance, contrast, and structural similarity between images. Its value ranges from 0 (completely dissimilar) to 1 (identical), formulated as:
$\operatorname{SSIM}(I,K)=\frac{\left(2\mu_{I}\mu_{K}+c_{1}\right)\left(2\sigma_{IK}+c_{2}\right)}{\left(\mu_{I}^{2}+\mu_{K}^{2}+c_{1}\right)\left(\sigma_{I}^{2}+\sigma_{K}^{2}+c_{2}\right)}$ (27)
where $\mu_{I}$, $\mu_{K}$ are the mean intensities; $\sigma_{I}$, $\sigma_{K}$ are the standard deviations; $\sigma_{IK}$ is the cross-covariance; and $c_{1}$, $c_{2}$are stabilization constants.
PSNR, derived from the logarithmic transformation of MSE, measures image distortion:
$\operatorname{PSNR}=10\cdot\log_{10}\left(\frac{\mathrm{MAX}_{I}^{2}}{\mathrm{MSE}}\right)$ (28)
where $\mathrm{MAX}_{I}$ denotes the maximum pixel value. Higher PSNR values indicate superior reconstruction quality.
4.3 Implementation Details
The code of our proposed USF-Net is built on the PyTorch framework with the Python programming language. The experimental platform and environment are as follows: a computer with Ubuntu 18.04, an Intel (R) Xeon (R) Gold 5318Y CPU @ 2.10 GHz, and two NVIDIA A40 GPUs with graphics memory of 48 GB. During the training phase, we set the initial learning rate to $10^{-3}$. To ensure the smooth training of the network, we used the SGD optimizer with 0.9 momentum and $10^{-4}$ weight decay. The training batch size is set to 4, and the learning rate will decay every 10 epochs until it reaches the minimum learning rate of $10^{-5}$. Consistent with the previous study [STANet], we employ the scheduled sampling strategy to mitigate generalization errors and enhance the ability to learn long-term spatiotemporal dynamics. Specifically, the occlusion rate $P$ applied to input images is progressively increased from 0 to 1 during the training duration. The proposed USF-Net is trained for 1,00 epochs on our dataset, and the training time was 4.8 h.
4.4 Results and Discussions
To evaluate the performance of our proposed method, we select several SOTA methods for comparison. These methods include both general spatiotemporal prediction methods (i.e., ConvLSTM [ConvLSTM], PredRNN [PredRNN], PredRNN++ [PredRNN++], MAU [MAU], LMC [LMC], and TAU [TAU]) and recent cloud image sequence extrapolation methods (i.e., CCLSTM [CCLSTM], CloudPredRNN++ [CloudPredRNN++], STANet [STANet], and MSTANet [MSTANet]). These DL methods represent different architectural paradigms (e.g., RNN-based, attention-based, hybrid models) and are widely recognized in the research community, allowing a comprehensive evaluation of USF-Net’s performance across multiple dimensions. All experimental results in this study are generated on our dataset by open-source codes.
Figure 8: Quantitative timestep-by-timestep comparison between our method and other methods on three metrics (a) MSE, (b) SSIM, and (c) PSNR. From 11 to 20 are the timesteps of extrapolation in order.
4.4.1 Quantitative Comparison
The quantitative evaluation results of our proposed USF-Net and other comparison methods on the ASI-CIS dataset are shown in Table 2. Among these tables, each row is the result for each method, and each column is the metric. The highest record is marked in bold. As demonstrated in Table 2, the proposed method achieves SOTA performance, attaining MSE, SSIM, and PSNR values of 37.18, 0.956, and 29.42, respectively, across all three evaluation metrics.
Table 2: Quantitative Comparison with Different Methods on the ASI-CIS Dataset. $\downarrow$ (or $\uparrow$) Indicates Lower (or Higher) is Better. The Best Results are Highlighted in Bold.
Method MSE($\downarrow$) SSIM($\uparrow$) PSNR($\uparrow$)
ConvLSTM (15NIPS)[ConvLSTM] 50.73 0.887 25.94
PredRNN (17NIPS)[PredRNN] 42.74 0.896 26.43
PredRNN++ (18ICML)[PredRNN++] 41.66 0.911 26.67
MAU (21NIPS)[MAU] 38.87 0.934 27.72
LMC (21CVPR)[LMC] 39.67 0.922 27.13
TAU (23CVPR)[TAU] 41.48 0.915 26.88
CCLSTM (21RS)[CCLSTM] 39.48 0.929 27.44
STANet (23TGRS)[STANet] 38.76 0.941 28.15
MSTANet (24TGRS)[MSTANet] 38.11 0.948 28.66
CloudPredRNN++ (25RS)[CloudPredRNN++] 38.44 0.945 28.34
USF-Net (Ours) 37.18 0.956 29.42
To further evaluate the long-term predictive capability of our method in cloud image sequence extrapolation tasks, we present the MSE, SSIM, and PSNR of each model at every timestep. As illustrated in Fig. 8, the proposed approach outperforms all baselines across metrics. The per-frame prediction curves of different models on the ASI-CIS dataset reveal distinct performance trends. Our method exhibits the weakest upward trajectory in MSE and the slowest decline in SSIM and PSNR, indicating superior stability over extended extrapolation horizons. Specifically, compared with the classic spatiotemporal sequence methods, our method introduces an SSM with a dynamic adaptive large-kernel selection mechanism, effectively addressing multi-scale variations in cloud imagery. When benchmarked against recent cloud extrapolation algorithms, our USF-Net exhibits a superior performance because the proposed UST can enhance the ability to integrate spatial and temporal features. By guiding spatial information refinement through temporal flow dynamics, the ability of our model to improve robust segmentation and capture long-term feature dependencies is enhanced. Consequently, our method achieves optimal performance even at the final timestep of extrapolation.
4.4.2 Qualitative Comparison
To further demonstrate the effectiveness of our method, we analyze the results of the comparison methods from a qualitative perspective. We selected some representative samples under diverse weather conditions, including sunny and cloudy/rainy scenarios, with all cloud imagery sequences exhibiting multi-scale cloud formations. Figs. 910 illustrate the visualization results for the representative samples. For cloud sequence extrapolation, both input and output sequences are configured with a length of 10 frames, captured at 30-second intervals. Specifically, we extracted the 1st and 6th frames (corresponding to timestamps $T$ = 1 and $T$ = 6) from each input sequence, while the 1st, 4th, 7th, and 10th output frames (corresponding to $T$ = 11, $T$ = 14, $T$ = 17, and $T$ = 20) are displayed. The first row contains the input and ground truth, and the remaining rows are the prediction results of each method.
Figure 9: Comparative extrapolation performance under sunny weather conditions is presented for ConvLSTM, PredRNN, PredRNN++, TAU, LMC, MAU, and our proposed method. All experiments are conducted on the ASI-CIS dataset, predicting the next ten images given the first ten observed frames.
Figure 10: Comparative extrapolation performance under cloudy/rainy weather conditions is presented for ConvLSTM, PredRNN, PredRNN++, TAU, LMC, MAU, and our proposed method. All experiments are conducted on the ASI-CIS dataset, predicting the next ten images given the first ten observed frames.
Compared with other methods, the proposed UTS-Net exhibits a more advanced prediction performance for cloud image sequences with different scales and deformations under diverse weather conditions. As shown in Fig. 9, conventional temporal networks omit boundary information during sequence extrapolation in sunny scenarios. Our method preserves complete contour and boundary details, benefiting from the spatial information branch incorporated in USF-Net that captures multi-scale cloud features. Moreover, the introduced TGM significantly enhances the capability to model long-range temporal dependency. Fig. 10 demonstrates that UTS-Net retains optimal prediction trajectories and cloud morphology even at the final extrapolation timestep. In addition, the DUM in UTS-Net effectively mitigates “ghosting effects” during sequence extrapolation as shown in Figs. 910. In summary, the proposed UTS-Net exhibits robust adaptability to multi-scale cloud extrapolation tasks across varying weather patterns while achieving SOTA performance in long-term spatiotemporal modeling.
Table 3: Complexity of Different Comparative Methods on the ASI-CIS Dataset. We Report the Parameters, Flops, Inference time, and MSE. $\downarrow$ Indicates Lower is Better. The Best Results are Highlighted in Bold.
Method Params(M) FLOPs(G) Inference time(ms) MSE($\downarrow$)
ConvLSTM [ConvLSTM] 18.0 215.3 17.3 50.73
PredRNN [PredRNN] 30.5 382.9 27.9 42.74
PredRNN++ [PredRNN++] 48.6 601.1 28.1 41.66
MAU [MAU] 19.3 281.1 16.8 38.72
LMC [LMC] 20.6 501.1 14.4 39.67
TAU [TAU] 44.7 294.4 19.7 41.48
CCLSTM [CCLSTM] 55.4 437.1 51.6 39.48
STANet [STANet] 26.5 462.9 16.8 38.76
MSTANet [MSTANet] 24.2 284.3 16.2 38.11
USF-Net (Ours) 23.8 266.4 15.8 37.18
4.4.3 Complexity Comparison
To evaluate the computation complexity of our method, we compare the parameter (Params(M)), floating-point operations (FLOPs), inference time and MSE of related methods on the ASI-CIS dataset. As shown in Table 3, the proposed USF-Net achieves optimal performance with the short inference time among all evaluated methods. While our method does not exhibit advantages in parameters and FLOPs compared to classic temporal prediction methods such as ConvLSTM and MAU, its performance gains fully justify the additional computational overhead. Furthermore, our method incurs lower computational costs than attention-based cloud extrapolation methods such as STANet and MSTANet due to the integration of the TGM. As evidenced by the inference time comparison, our method achieves near-optimal efficiency (second only to LMC), which is sufficient for cloud imagery captured at 30-second intervals and aligns with the requirements of ultra-short-term PV power forecasting. Therefore, our proposed UTS-Net establishes a better accuracy-speed trade-off in ground-based remote sensing cloud image sequence extrapolation.
4.5 Ablation Study
To further verify the effectiveness of the different modules of our proposed UTS-Net, we also conducted a comprehensive ablation study. The proposed UTS-Net employs an encoder-decoder framework with a unified spatiotemporal module (comprising SSM-based spatial branch, TAM-based temporal branch, and TGM-based dynamic spatiotemporal module) and a decoder structure incorporating DUM. Therefore, we conduct different experiments to verify the proposed modules on the ASI-CIS dataset. First, we select UTS-Net as the baseline. Then, we incrementally remove the SSM, TAM and TGM from the baseline to verify the effectiveness of the proposed USTM. Finally, we remove the DUM from the baseline to verify its validity.
[…]
Table 4: Ablation Experimental Results on the ASI-CIS Dataset. $\downarrow$ (or $\uparrow$) Indicates Lower (or Higher) is Better. The Best Results are Highlighted in Bold.
Version SSM TAM TGM DUM MSE($\downarrow$) SSIM($\uparrow$) Params(M)
Baseline 37.18 0.956 23.8
1 39.96 0.918 23.3
2 41.74 0.906 22.9
3 40.24 0.915 23.1
4 38.65 0.942 23.5
5 SA 37.19 0.951 24.6
Table 5: Ablation Experimental Results of the Number of Decomposed Large Kernels with the RF being 23.
RF (k,d) Sequence Number Inference time(ms) MSE($\downarrow$)
23 (23, 1) 1 16.6 38.21
[…]
Table 6: Ablation Experimental Results with Different RFs of the Dynamic Large-kernel Selection. RF = 23 Corresponds to Our Proposed Method.
RF (k,d) Sequence Inference time(ms) Params(M) MSE($\downarrow$)
[…]
Figure 11: To highlight the impact of the GAU, representative samples are presented: (a) ground truth data, (b) prediction from USF-Net without the DUM, and (c) prediction from USF-Net with DUM. Regions marked by red boxes indicate areas with prediction deficiencies in the absence of DUM, while green boxes demonstrate substantial improvements achieved through DUM integration.
Moreover, to evaluate the impact of dynamic large-kernel selection on cloud image sequence extrapolation performance, the ablation study is conducted on the selection of the multi-scale large-kernel in the spatial branch. When the RF is fixed at 23, we conduct an experiment on the number of large kernel decompositions. The experimental results, as shown in Table 5, achieve a good trade-off between speed and accuracy by decomposing the large kernel into two depth-wise kernels, resulting in excellent performance in both inference time and MSE. In addition, we configured the RF as 11, 21, 23, 29, and 39, where RF = 23 corresponds to our proposed method. As shown in Table 6, decomposing large kernels into two depth-wise components effectively captures multi-scale cloud motion patterns, significantly improving prediction accuracy for cloud image sequences. However, excessively small or large RFs can hinder the performance of the USF-Net. The performance degrades when the RF exceeds 23 due to excessive detail loss when decomposed kernels encounter smaller-scale cloud structures. The experimental results demonstrate that our selected large kernel decomposition strategy achieves an optimal balance between prediction performance and computational efficiency.
5 Conclusion
Accurate and efficient extrapolation of ground-based cloud image sequences is a critical enabling technology for the stable integration of photovoltaic power systems. In this paper, we addressed key limitations in existing deep learning methods by proposing USF-Net, a novel framework for cloud image sequence extrapolation that introduces spatiotemporal architecture to unify the modeling of spatial and temporal information a novel framework that unifies the modeling of spatial and temporal information. Our contributions are threefold. First, we introduced a unified spatiotemporal architecture where temporal flow information explicitly guides spatial feature learning via a novel Temporal Guidance Module (TGM). Second, we designed a Unified Spatiotemoral Module (USTM) that contains a Spatial Selection Module (SSM) to dynamically capture multi-scale cloud context and a Temporal Agent Attention Module (TAM) to model long-range temporal dependencies with linear complexity efficiently. Third, we developed a Dynamic Update Module (DUM) in the decoder that leverages initial temporal states to effectively mitigate the “ghosting effect” and preserve motion fidelity. Extensive experiments on our newly proposed high-resolution ASI-CIS dataset demonstrate that USF-Net significantly outperforms existing SOTA methods. Furthermore, ablation studies rigorously validate the effectiveness of each of our proposed modules. The results confirm that USF-Net establishes a new benchmark for the task, achieving superior prediction accuracy while maintaining high computational efficiency. Future work will focus on extending USF-Net to additional photovoltaic power prediction tasks, further enhancing the method’s real-time inference capability.
CRediT authorship contribution statement
Penghui Niu: Writing - original draft, Writing – review &\& editing, Visualization, Validation, Software, Methodology, Investigation, Funding acquisition. Taotao Cai: Writing – review &\& editing, Methodology, Investigation, Formal analysis. Jiashuai She: Writing - original draft, Writing – review &\& editing, Validation, Data curation. Yajuan Zhang: Resources, Project administration, Investigation. Junhua Gu: Supervision, Resources, Project administration, Funding acquisition, Data curation. Ping Zhang: Resources, Project administration, Funding acquisition. Jungong Han: Writing – review &\& editing, Supervision, Methodology. Jianxin Li: Writing – review &\& editing, Supervision, Project administration, Methodology.
Declaration of competing interest
The authors declare that they have no known competing finan cial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The proposed ASI-CIS dataset and source code will be available at https://github.com/she1110/ASI-CIS.
Acknowledgments
The research was supported by the National Nature Science Foundation of China under Grant No. 62206085, supported by the Innovation Capacity Improvement Plan Project of Hebei Province under 22567603H, and supported by the Interdisciplinary Postgraduate Training Program of Hebei University of Technology under HEBUT-Y-XKJC-2022101.